OpenStudy (lena772):

Functions f(x) and g(x) are shown below. f(x) = x^2 g(x) = x^2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? Left by 4 units Right by 4 units Left by 8 units Right by 8 units I'm pretty sure it goes to the right, but I don't know by how many units.

OpenStudy (solomonzelman):

\[ x^2 + 8x + 16=(x+4)^2\]

hero (hero):

Hint: Rewrite g(x) in the form (x + a)^2

OpenStudy (lena772):

So to the right by 4 units?

OpenStudy (lena772):

why does it go left if it is increasing

hero (hero):

Because what you see is an illusion: (x + 4)^2 = (x - (-4))^2 You re-write the expression in the form on the RHS to see which way the graph has shifted. -4 means "4 units to the left"